### v1

# Markovian Quantum Computation

Abstract
We propose a general prototype of quantum computation based on the finding of many-body ground states. The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir. Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system towards its ground state. At the same time, measurement of the discarded qubits hints at the energy level structure of many-body systems. We show that quantum computation based on this Markovian process is equivalent in its computational power to the one based on quantum circuits. We exemplify the scheme with several illustrative use cases for combinatorial optimization problems, discussing the efficiency and the issue of local energy minima. We also discuss its application to the preparation of quantum many-body ground states for gapped systems.
I Introduction
Computing many-body ground states is extremely hard in general: for classical Hamiltonians, it falls into the NP-hard complexity class, and for quantum Hamiltonians, it is classified as QMA-hard [1, 2, 3]. The difficulty arises mainly from the geometric frustration, i.e., the stark mismatch between local and global energy minima. The problem is combinatorial in nature: one needs to find the optimal configurations of data among exponentially many, mutually conflicting ones. This is indeed the core challenge of combinatorial optimization problems [4]. It is well-established that combinatorial optimization problems can be reduced to the task of finding the ground states of Ising-like spin models [5, 6, 7].
It is widely believed that quantum computers are not universally efficient in solving general ground-state problems, even for classical Hamiltonians. Nonetheless, we still need to find out specific classes of such problems that benefit from quantum computation and are yet practically useful. Also, there’s a need to devise diverse quantum algorithms for them, each having its unique advantages. We have a number of tools available for such purposes [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] (Ref. [21] contains a nice summary of earlier schemes). The most prominent is adiabatic quantum computation (AQC) [10, 11, 12]. The AQC exploits the adiabatic evolution of the instantaneous ground state when the Hamiltonian is adiabatically varied. The time scale of the operation is thus governed by the minimal spectral gap during the evolution. However, analyzing the spectral gap is extremely difficult, leaving the performance of AQC largely unknown. There are also hybrid approaches, such as quantum approximate optimization algorithms (QAOA) [14, 15, 13] and variational quantum eigensolvers (VQE) [16, 17, 18], aided by classical optimization subroutines.
As mentioned, the difficulty in analyzing AQC arises as the initial and final Hamiltonians intermingle. On the other hand, the properties of the final Hamiltonian alone can be largely unveiled, even though its ground state remains unknown. Given this, it is natural to consider schemes to directly cool the final Hamiltonian system to its ground state. While earlier schemes based on dissipative engineering [19, 20, 21] appear to achieve this goal, a closer inspection reveals that they actually lack desirable properties as an ideal cooling method. To clarify this, it is instructive to discuss classical and quantum Hamiltonians separately. By the former, we mean those Hamiltonians diagonalized in the computational basis.
For classical Hamiltonians, the earlier dissipative schemes do not provide any quantum advantage. There are a few fundamental reasons. They implement a completely-positive trace-preserving (CPTP) map that transforms the state $\rho$ to $\sum_{k}E_{k}\rho E_{k}^{\dagger}$ [22]. Here, each Kraus operator $E_{k}$ directly accesses the associated local part of the system, driving it into a certain direction. However, as the energy of the whole system is a global quantity, determining the optimal direction based solely on a small part is not possible (note that this was the very origin of the geometric frustration, rendering the problem intractable). For this reason, the state is randomized when the measurement projects the state into an undesired subspace. Note, however, that this entire process is essentially equivalent to a mere random guess. Furthermore, for classical Hamiltonians, they do not involve any operations in non-computational bases. This means that they are, in fact, identical to classical probabilistic models. In the case of quantum Hamiltonians, the problem is more involving: one need to find an appropriate basis containing the ground state. This is the place where the above CPTP map comes into play. Apart from the role of fixing the basis, however, the above mentioned limitations persist.
An optimal cooling should involve quantum transitions that take place exclusively in the direction of lowering energy. Here, a mechanism to prevent heating can be ensured only by an external agent, which globally accesses the system, gathers its energy, and relaxes it into a zero-temperature Markovian reservoir. This process is exactly given by cavity cooling methods [23]. The aim of this paper is to materialize this concept as a general quantum computation model and discuss its efficiency in some illustrative scenarios. Our model can be paralleled with AQC for their similarity in motivation. On the other hand, unlike AQC that is a unitary process, our model exploits a Markovian process. In order to stress this and for the sake of brevity, we will refer our model to Markovian quantum computation (MQC) hereafter.
Being a Markovian process, MQC inherits the intrinsic advantages of the earlier dissipative schemes, alleviating the demand of precise system controls to some extent. On top of this, MQC takes true quantum advantages. In fact, we show that MQC is equivalent to quantum circuit models in terms of computational power. This argument is similar to the established one regarding the computational equivalence between AQC and quantum circuit models [11]. All three frameworks are thus computationally equivalent. However, this assertion needs to be interpreted with care. For instance, the equivalence of AQC to MQC is established by translating the unitary evolution of AQC into a quantum circuit, which is then converted into a Hamiltonian for MQC. However, this resulting Hamiltonian differs from the original problem Hamiltonian of the AQC. It is unclear if the feasibility of ground-state preparation by AQC implies that of the same Hamiltonian by MQC, and vice versa.
Any ground-state preparation scheme, including AQC, inevitably suffers from the issue of local energy minima [6, 24, 25]. In MQC, one can design tunneling transitions out of local minima if the information is given on their nature. This flexibility also constitutes the advantage of MQC. In practice, one could try various transition terms in the Hamiltonian, while observing the occurrence of the transition from the detection of cavity photons. This, in turn, provides a method to inspect the energy-level structure of the Hamiltonian.
In this work, our main focus will be placed on MQC for combinatorial optimization problems. Such problems involve classical Hamiltonians and have a number of favorable features for MQC. We consider two extreme cases. The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size. This case turns out to be equivalent to Grover’s quantum search algorithm. The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begins to pose challenges. In this case, high-order transitions, required to get out of local energy minima, become the bottleneck. These transitions require a time that is exponentially large in the order, which can increase with the system size in general. Finally, we briefly discuss the case of quantum Hamiltonians. This case is heavily system-dependent and harder to analyze. We argue that the spectral gap above the ground state determines the basic time scale of the cooling.
II Notation and working principles
Consider an $N$-qubit system described by Hamiltonian
$H_{P}=\sum_{z=0}^{2^{N}-1}E(z)|z\rangle_{P}\langle z|,$ (1)
where $E(z)\leq E(z^{\prime})$ for $z<z^{\prime}$. Let us call $H_{P}$ a problem Hamiltonian. Our aim is to find the ground state $|0\rangle_{P}$. Transitions among the energy levels $|z\rangle_{P}$ are allowed by introducing in the Hamiltonian what we call a transition term $H_{T}$, which does not commute with $H_{P}$. $H_{T}$ should be designed to suit the need of the given problem. We will consider two examples below. Here, we assume the energy level spacings of $H_{P}$ are characterized by an energy scale $\Delta>0$, and so is the norm $\|H_{T}\|$. Our aim is to introduce a parameter $0<\lambda\ll\Delta$ to treat $\lambda H_{T}$ as a perturbation to $H_{P}$.
If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, ignoring the influence of the degeneracy. In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition. Let us denote by $|\cdot\rangle_{m}$ the photon number state of the $m$-th cavity. If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to avoid heating.
Introducing $M$ different cavity modes, our final Hamiltonian reads
$\begin{split}H&=H_{P}+\sum_{m=1}^{M}\omega_{m}c_{m}^{\dagger}c_{m}\\ &\quad+\lambda H_{T}\otimes\left[a_{0}I+\sum_{m=1}^{M}(c_{m}+c_{m}^{\dagger})\right].\end{split}$ (2)
Here, $a_{0}\in\{0,1\}$ determines whether the transitions preserving the cavity photon numbers are allowed. This term helps the state escaping from local minima through high-order transitions. In practice, it is generally enough to restrict the maximum number of photons in each cavity to one. In this case, $M$ cavity modes can be simulated with $M$ qubits.
Initially, the system is prepared in state $2^{-N/2}\sum_{z=0}^{2^{N}-1}|z\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$. A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state. Here, the measurement of the cavity photons is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
The computation is done by repeating this cycle. The required number of cycles depends on the problem and strategy. One possible strategy is to repeat until the cavity photon is not detected for an enough number of cycles. This signals with high probability that the system has reached either the ground state or local minima. For NP problems that are guaranteed to have at least one solution, deciding which is the case is straightforward. Otherwise, one may repeat the same computation to increase the confidence. Given the unknown ground state not being verifiable in general, the decision is probabilistic after all.
III Equivalence to circuit models
Figure 1: The transition diagram into which a quantum circuit is converted for MQC. $|\phi_{t}\rangle_{P}$, $|t\rangle_{C}$, and $|\cdot\rangle_{1}$ denotes, respectively, the state of the quantum circuit at the $t$-th time step, the clock state to label the time step, and the number state of the cavity. $\Delta$ is the energy difference between the adjacent levels and $\lambda$ is the Rabi frequency of the designated transition. The wavy lines represent the transition after detecting and clearing the cavity photon.
Ref. [11] shows that AQC is equivalent to the quantum circuit model in computational power up to polynomial overhead. The forward direction of the equivalence is rather straightforward because the entire procedure of AQC is a unitary transformation, which can be simulated efficiently in quantum circuits. Showing the other direction replies on the clock state $|t\rangle_{C}$ that is tagged to the state at the $t$-th time step of the quantum-circuit computation.
A similar equivalence relation can be shown for MQC. Again, showing that MQC can be simulated efficiently in quantum circuits is straightforward. We thus focus on the other direction. Suppose a certain quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step. Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$. The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition. This is enabled by choosing the terms in Hamiltonian (2) as
$H_{P}$ $=-\sum_{t=0}^{T}t\Delta|t\rangle_{C}\langle t|,$ (3)
[…]
where the parameter $\Delta>0$ is an arbitrary energy scale, and the other parmeters are chosen as $M=1$, $\omega_{1}=\Delta$, and $\alpha_{0}=0$. If the initial state is chosen to be $|\phi_{0}\rangle_{P}|0\rangle_{C}|0\rangle_{1}$, the entire transition dynamics is restricted to a subspace spanned by $\{|\phi_{t}\rangle_{P}|t\rangle_{C},0\leq t\leq T\}$, as depicted in Fig. 1. It can be seen that the average time needed to reach the ground state, which is the final state of the quantum circuit, is polynomial in $T$. As the transition rate $\lambda$ is restricted by the energy gap $\Delta$, the computation time is $\mathcal{O}(T/\Delta)$.
IV Combinatorial optimization by Markovian quantum computation
IV.1 Notation
Consider combinatorial optimization problems that aim to find out the optimal solutions from a set of $2^{N}$ possible configurations represented by $N$-bit string $z\equiv z_{1}z_{2}\cdots z_{N}$. The optimality of each configuration is evaluated by a non-negative cost function $E(z)$. The objective is to identify configurations $z$ that minimize $E(z)$. This problem is straightforwardly translated into the ground-state finding problem, where the cost function $E(z)$ is identified with the energy eigenvalue in Hamiltonian (1). Here, we set $E(z)$ to be integer multiples of a parameter $\Delta>0$. The cavity frequencies in Hamiltonian (2) can then be chosen as integer multiples of $\Delta$. By choosing $M$ to be the maximum of $E(z)$, all possible transitions in the problem Hamiltonian can be coupled to a cavity transition. This imposes a polynomial overhead to the computational resource.
IV.2 Reproduction of Grover’s search algorithm
Grover’s quantum search algorithm is paradigmatic in quantum information theory [26, 22]. It can be recast as an algorithm to find the zeros of a binary function $f(z)\in\{0,1\}$ from the set of $2^{N}$ configurations $z$. Provided that $f(z)$ is computed only by an external agent, called the oracle, Grover’s algorithm requires $\mathcal{O}(2^{N/2})$ oracle calls to find the solution, while the number becomes $\mathcal{O}(2^{N})$ for the best classical algorithm. The quadratic speedup of Grover’s algorithm is proven to be optimal unless the structure of the function $f(z)$ is exploited somehow [27, 28, 29].
To perform an analogous task in MQC, we adjust the Hamiltonian (2) as
$\begin{split}H&=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}\\ &\quad+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),\end{split}$ (5)
where $I_{i}$ and $X_{i}$ denote the identity and Pauli $X$ operators acting on qubit $i$, respectively. Let $n_{0}$ be the number of zeros of $f(z)$. For the initial state $2^{-N/2}\sum_{z}|z\rangle_{P}|0\rangle_{1}$, the problem Hamiltonian $H_{P}$ effectively becomes two-dimensional, where the Hilbert space is spanned by
$|\phi_{0}\rangle_{P}=\frac{1}{\sqrt{n_{0}}}\sum_{f(z)=0}|z\rangle_{P}$ (6)
and
$|\phi_{1}\rangle_{P}=\frac{1}{\sqrt{2^{N}-n_{0}}}\sum_{f(z)\not=0}|z\rangle_{P}.$ (7)
In this basis, the off-diagonal element becomes ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$. The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm. One advantage of MQC implementation is that the end of computation is signaled by a detection of the cavity photon. This is especially beneficial when the number of solutions $n_{0}$ is unknown.
In this example, the problem Hamiltonian has only two energy levels corresponding to the solution states and the rest. A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels. It can be seen that with the transition term chosen above, more energy levels just make the computation slower. To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level. The increase in the number of energy levels leads to the decrease in the degeneracy of each level, which diminishes the collective effect. Moreover, it results in an increased number of transitions required to reach the ground state.
IV.3 Alternative approach
In the previous example reproducing Grover’s search algorithm, the transition term generates an all-to-all interaction with the identical transition strength. This makes the problem Hamiltonian completely featureless except for each energy level having a different degeneracy. While the absence of any local energy minimum is advantageous, the transition strength decreases superpolynomially with the number of qubits.
In this subsection, we consider an opposite limit, where the transition term is given by
$H_{T}=\sum_{i=1}^{N}X_{i}.$ (8)
The off-diagonal element ${}_{P}\langle z|H_{T}|z^{\prime}\rangle_{P}$ is then non-vanishing and equals one, independently of $N$, only when the Hamming distance between $z$ and $z^{\prime}$ is one. While the transition strength is maximum, local energy minima now become an issue. As the distance between two configurations is at most $N$, any state $|z\rangle_{P}$ can reach the ground state in at most $N$ transitions in principle. However, we do not know a general rule to take the right path for every state. For this reason, we take the direction of lowering the energy, even though it does not coincide with the direction of approaching the ground state. This mismatch is the origin of local energy minima.
Note that every state $|z\rangle_{P}$ is linked to $N$ different states by transition. It is instructive to envisage this as an $N$-regular graph with $2^{N}$ vertices. Each vertex is endows with a potential determined by the problem Hamiltonian, and the transition occurs along the edges toward the direction of not increasing the potential. Once the population is trapped in a local potential minimum, a high-order transition is needed to get out of it.
To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\rightarrow|z_{1}\rangle_{P}\rightarrow\cdots|z_{n}\rangle_{P}$. For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved. For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$. The perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$. However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition. Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$. Now taking the cavity modes into account, the last condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$. This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists. By incorporating multiple cavity modes, various tunnelling-out transition channels can appear. For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play a crucial role in overcoming local minima.
To illustrated the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [30, 22]). To be specific, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem. The problem Hamiltonian is constructed as follows. Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$. For a given value of $\bar{z}$, we turn each step of the calculation into an energy term in the problem Hamiltonian. Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian. For example, the first energy term has value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{b}$ and $\Delta$ otherwise, and so on. The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.
Figure 2: Numerical simulation of MQC for integer factoring of 35. Dotted curves represent the energy with respect to the number of cooling cycles, averaged over $10^{3}$ samples, for (a) $\alpha_{0}=1$ and (b) $\alpha_{0}=0$. Gray curves represent typical individual trajectories.
Fig. 2 shows the results of our numerical simulation for $\bar{z}=35$. We take the transition term as in Eq. (8) and set the parameters in Hamiltonian (2) as $M=3$, $\omega_{m}=m\Delta$, $\lambda=\Delta/10$, and $a_{0}=1$. The duration of a single cooling cycle is chosen to be $\pi/2\lambda$. The simulation is performed by exact diagonalization. Fig. 2(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state. As a comparison, we have changed $a_{0}$ to zero in Fig. 2(b) to test the effectiveness of the aforementioned mechanism for tunneling out from local minima. In the absence of such a mechanism, the system almost always evolves into local minima, which appear to be present in the first excited level.
A remark is in order. Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears to be effective, as demonstrated in Fig. 2(a). However, it should be noted that the optimal choice of $\lambda$ depends on the characteristics and size of the problem. In general, smaller $\lambda$ is preferable as the collective effect enhance the transition rate. This effect is particularly important for combinatorial optimization problems, which are characterized by large degeneracies in the problem Hamiltonian. As the degeneracies of low energy levels are much smaller in general, the collective effect is also weaker in the low-energy part of the spectrum. This means that there exists a parametric regime of $\lambda$ where the heating is sufficiently avoided for the ground state, but not for higher energy states. On the other hand, by choosing smaller $\lambda$, leading to slower transitions, heating could be avoided throughout the entire energy spectrum. While the cooling works in both cases, it is unclear which one is more efficient.
V Discussion
In this work, we have introduced the notion of Markovian quantum computation, along with a few illustrative examples. The idea is quite straightforward: cooling the system with a problem Hamiltonian through energy-lowering transitions, as in cavity cooling. An essential part here is to choose an appropriate transition term that generates off-diagonal elements in the diagonal basis of the problem Hamiltonian. There exists a trade-off between the connectivity and strength of the transitions. We have discussed two opposite limits when the ground state is the solution of a combinatorial optimization problem. In both cases, the running time generally increases superpolynomially with the system size. It is unclear if the performance can be substantially improved by interpolating between the two opposites.
MQC can also be used to find the ground state of general quantum Hamiltonians. The most important requirement is that the ground state is not heated during the process to make it a steady state. This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap can be resolved by a proper choice of the transition term and the parameters. Consequently, the inverse of the spectral gap determines the basic time scale of the computation. In principle, a finite number of cavity modes can cover the transitions across the entire spectrum of the problem Hamiltonian, where the required number of cavity modes is given roughly by the ratio of the energy range of the problem Hamiltonian to the characteristic transition rate. This means that the number of cavity modes increases polynomially with the system size for gapped Hamiltonians. However, choosing an appropriate transition term and estimating the resulting cooling time depend heavily on the system and are analytically challenging.
Acknowledgements.
This work was supported by the National Research Foundation (NRF) of Korea under Grant No. NRF-2022R1A4A1030660.

### v2

@@
-Markovian Quantum Computation
+Quantum Computation by Cooling
 Abstract
-We propose a general prototype of quantum computation based on the finding of many-body ground states.
+Adiabatic quantum computation is a paradigmatic model aiming to solve a computational problem by finding the many-body ground state encapsulating the solution.
+However, its use of an adiabatic evolution depending on the spectral gap of an intricate many-body Hamiltonian makes its analysis daunting.
+While it is plausible to directly cool the final gapped system of the adiabatic evolution instead, an analysis of such a scheme is surprisingly missing.
+Here, we propose a specific Hamiltonian model for this purpose.
 The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir.
-Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system towards its ground state.
-At the same time, measurement of the discarded qubits hints at the energy level structure of many-body systems.
-We show that quantum computation based on this Markovian process is equivalent in its computational power to the one based on quantum circuits.
-We exemplify the scheme with several illustrative use cases for combinatorial optimization problems, discussing the efficiency and the issue of local energy minima.
-We also discuss its application to the preparation of quantum many-body ground states for gapped systems.
+Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system toward its ground state.
+At the same time, the measurement of the discarded qubits hints at the energy level structure of the system as a return.
+We show that quantum computation based on this cooling procedure is equivalent in its computational power to the one based on quantum circuits.
+We then exemplify the scheme with a few illustrative use cases for combinatorial optimization problems.
+In the first example, the cooling is free from any local energy minima, reducing the scheme to Grover’s search algorithm with a few improvements.
+In the second example, the cooling suffers from abundant local energy minima.
+To circumvent this, we implant a mechanism in the Hamiltonian so that the population trapped in the local minima can tunnel out by high-order transitions.
+We support this idea with a numerical simulation for a particular combinatorial optimization problem.
+We also discuss its application to preparing quantum many-body ground states, arguing that the spectral gap is a crucial factor in determining the time scale of the cooling.
 I Introduction
+Identifying the ground states of quantum many-body Hamiltonians and understanding their characteristics are fundamental in physics.
+The ground states encode the spectral properties of the Hamiltonian [1, 2, 3, 4, 5, 6, 7] and represent the essential quantumness of the system.
+The latter becomes particularly evident in gapped systems, in which the ground state predominantly determines the entire properties at sufficiently low temperature.
+This lays the foundation for various exotic quantum phenomena, such as quantum Hall effects and superconductivity [8].
 Computing many-body ground states is extremely hard in general: for classical Hamiltonians, it falls into the NP-hard complexity class, and for quantum Hamiltonians, it is classified as QMA-hard [9, 10, 11].
@@
 The most prominent is adiabatic quantum computation (AQC) [18, 19, 20].
-The AQC exploits the adiabatic evolution of the instantaneous ground state when the Hamiltonian is adiabatically varied.
+The AQC exploits the adiabatic evolution of the instantaneous ground state when the Hamiltonian is varied.
 The time scale of the operation is thus governed by the minimal spectral gap during the evolution.
@@
 Given this, it is natural to consider schemes to directly cool the final Hamiltonian system to its ground state.
-While earlier schemes based on dissipative engineering [19, 20, 21] appear to achieve this goal, a closer inspection reveals that they actually lack desirable properties as an ideal cooling method.
+While earlier schemes based on dissipative engineering [27, 28, 29] seem to achieve this goal, a closer inspection reveals that they actually lack desirable properties as a cooling method.
 To clarify this, it is instructive to discuss classical and quantum Hamiltonians separately.
 By the former, we mean those Hamiltonians diagonalized in the computational basis.
-For classical Hamiltonians, the earlier dissipative schemes do not provide any quantum advantage.
+For classical Hamiltonians, the earlier dissipative schemes are reduced to a classical random search.
 There are a few fundamental reasons.
@@
 This means that they are, in fact, identical to classical probabilistic models.
-In the case of quantum Hamiltonians, the problem is more involving: one need to find an appropriate basis containing the ground state.
+In the case of quantum Hamiltonians, the problem is more involving: one needs to find an appropriate basis containing the ground state.
 This is the place where the above CPTP map comes into play.
-Apart from the role of fixing the basis, however, the above mentioned limitations persist.
-An optimal cooling should involve quantum transitions that take place exclusively in the direction of lowering energy.
-Here, a mechanism to prevent heating can be ensured only by an external agent, which globally accesses the system, gathers its energy, and relaxes it into a zero-temperature Markovian reservoir.
-This process is exactly given by cavity cooling methods [23].
-The aim of this paper is to materialize this concept as a general quantum computation model and discuss its efficiency in some illustrative scenarios.
+Apart from the role of fixing the basis, however, the above-mentioned limitations persist.
+Simply put, both cooling and heating coexist in the above-mentioned schemes.
+A genuine cooling scheme should rely on quantum transitions that take place exclusively in the direction of lowering energy.
+It is important to understand that a heating-prevention mechanism requires an external agent, which globally accesses the system, gathers its energy, and relaxes it into a Markovian reservoir.
+This process is exactly provided by cavity cooling methods [31].
+This paper aims at materializing this concept as a general quantum computation model and discuss its efficiency in some illustrative scenarios.
 Our model can be paralleled with AQC for their similarity in motivation.
-On the other hand, unlike AQC that is a unitary process, our model exploits a Markovian process.
-In order to stress this and for the sake of brevity, we will refer our model to Markovian quantum computation (MQC) hereafter.
-Being a Markovian process, MQC inherits the intrinsic advantages of the earlier dissipative schemes, alleviating the demand of precise system controls to some extent.
-On top of this, MQC takes true quantum advantages.
-In fact, we show that MQC is equivalent to quantum circuit models in terms of computational power.
+On the other hand, unlike AQC, our model exploits a non-unitary process.
+For brevity, we will refer to our model as cooling-based quantum computation (CQC) hereafter.
+Being a dissipative process, CQC inherits the intrinsic advantages of the earlier dissipative schemes, alleviating the demand of precise system controls to some extent.
+On top of this, CQC takes true quantum advantages.
+In fact, we show that CQC is equivalent to quantum circuit models in terms of computational power.
 This argument is similar to the established one regarding the computational equivalence between AQC and quantum circuit models [19].
@@
 However, this assertion needs to be interpreted with care.
-For instance, the equivalence of AQC to MQC is established by translating the unitary evolution of AQC into a quantum circuit, which is then converted into a Hamiltonian for MQC.
+For instance, the equivalence of AQC to CQC is established by translating the unitary evolution of AQC into a quantum circuit, which is then converted into a Hamiltonian for CQC.
 However, this resulting Hamiltonian differs from the original problem Hamiltonian of the AQC.
-It is unclear if the feasibility of ground-state preparation by AQC implies that of the same Hamiltonian by MQC, and vice versa.
+It is an open question whether the feasibility of ground-state preparation by AQC implies that of the same Hamiltonian by CQC, and vice versa.
 Any ground-state preparation scheme, including AQC, inevitably suffers from the issue of local energy minima [14, 32, 33].
-In MQC, one can design tunneling transitions out of local minima if the information is given on their nature.
-This flexibility also constitutes the advantage of MQC.
-In practice, one could try various transition terms in the Hamiltonian, while observing the occurrence of the transition from the detection of cavity photons.
+In CQC, one can design tunneling transitions out of local minima if the information is given on their nature.
+This flexibility also constitutes the advantage of CQC.
+In practice, one could try various transition terms in the Hamiltonian, while observing the occurrence of the transition by monitoring cavity photons.
 This, in turn, provides a method to inspect the energy-level structure of the Hamiltonian.
-In this work, our main focus will be placed on MQC for combinatorial optimization problems.
-Such problems involve classical Hamiltonians and have a number of favorable features for MQC.
+For example, the presence or absence of a finite spectral gap above the ground state could be inspected for unsolvable Hamiltonians.
+In this work, much of our focus will be placed on CQC for combinatorial optimization problems, which have favorable features for CQC and allow for the analytical treatment.
 We consider two extreme cases.
 The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size.
-This case turns out to be equivalent to Grover’s quantum search algorithm.
-The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begins to pose challenges.
-In this case, high-order transitions, required to get out of local energy minima, become the bottleneck.
-These transitions require a time that is exponentially large in the order, which can increase with the system size in general.
+This case turns out to be equivalent to Grover’s quantum search algorithm, aside from some additional advantages that CQC brings.
+The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begin to pose challenges.
+In this case, high-order transitions, which are required to get out of local energy minima, become the bottleneck.
+These transitions require a time exponentially large in the order, which can increase with the system size in general.
+We perform the numerical simulation for a particular combinatorial optimization problem to support this idea.
 Finally, we briefly discuss the case of quantum Hamiltonians.
-This case is heavily system-dependent and harder to analyze.
-We argue that the spectral gap above the ground state determines the basic time scale of the cooling.
+This case is heavily system-dependent and more complicated to analyze.
+We argue that the spectral gap above the ground state is a crucial factor in determining the time scale of the cooling.
+This is a natural consequence of the energy-time uncertainty principle.
 II Notation and working principles
@@
 Let us call $H_{P}$ a problem Hamiltonian.
-Our aim is to find the ground state $|0\rangle_{P}$.
+We aim to find the ground state $|0\rangle_{P}$.
 Transitions among the energy levels $|z\rangle_{P}$ are allowed by introducing in the Hamiltonian what we call a transition term $H_{T}$, which does not commute with $H_{P}$.
-$H_{T}$ should be designed to suit the need of the given problem.
-We will consider two examples below.
+$H_{T}$ should be designed to suit the needs of the given problem.
+We will consider two simple examples below.
 Here, we assume the energy level spacings of $H_{P}$ are characterized by an energy scale $\Delta>0$, and so is the norm $\|H_{T}\|$.
 Our aim is to introduce a parameter $0<\lambda\ll\Delta$ to treat $\lambda H_{T}$ as a perturbation to $H_{P}$.
-If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, ignoring the influence of the degeneracy.
+If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, where the influence of the degeneracy is ignored.
 In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition.
-Let us denote by $|\cdot\rangle_{m}$ the photon number state of the $m$-th cavity.
-If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to avoid heating.
+Here, the subscript $m$ is the index for different cavities.
+Let us denote by $|\cdot\rangle_{m}$ the photon number state of the cavity.
+If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to prevent heating.
 Introducing $M$ different cavity modes, our final Hamiltonian reads
@@
 Here, $a_{0}\in\{0,1\}$ determines whether the transitions preserving the cavity photon numbers are allowed.
-This term helps the state escaping from local minima through high-order transitions.
+This term helps the state escape from local minima through high-order transitions.
 In practice, it is generally enough to restrict the maximum number of photons in each cavity to one.
@@
 A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state.
-Here, the measurement of the cavity photons is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
+Here, the measurement of the cavity photon is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
 The computation is done by repeating this cycle.
@@
 Otherwise, one may repeat the same computation to increase the confidence.
-Given the unknown ground state not being verifiable in general, the decision is probabilistic after all.
+Given that the unknown ground state is not verifiable in general, the decision is probabilistic after all.
+We will implicitly assume that the evolution by the Hamiltonian (2) is simulated on a universal quantum computer, which is reasonable given the complexity of the Hamiltonian.
+In our analysis below, the polynomial overhead associated with this implementation is not taken into account.
 III Equivalence to circuit models
-Figure 1: The transition diagram into which a quantum circuit is converted for MQC.
+Figure 1: The transition diagram into which a quantum circuit is converted for CQC.
 $|\phi_{t}\rangle_{P}$, $|t\rangle_{C}$, and $|\cdot\rangle_{1}$ denotes, respectively, the state of the quantum circuit at the $t$-th time step, the clock state to label the time step, and the number state of the cavity.
@@
 [19] shows that AQC is equivalent to the quantum circuit model in computational power up to polynomial overhead.
-The forward direction of the equivalence is rather straightforward because the entire procedure of AQC is a unitary transformation, which can be simulated efficiently in quantum circuits.
-Showing the other direction replies on the clock state $|t\rangle_{C}$ that is tagged to the state at the $t$-th time step of the quantum-circuit computation.
-A similar equivalence relation can be shown for MQC.
-Again, showing that MQC can be simulated efficiently in quantum circuits is straightforward.
+The forward direction of the equivalence is straightforward because the entire procedure of AQC is a unitary transformation, which can be simulated efficiently in quantum circuits.
+Showing the other direction relies on the clock state $|t\rangle_{C}$ tagged to the state at the $t$-th time step of the quantum-circuit computation.
+A similar equivalence relation can be shown for CQC.
+Again, showing that CQC can be simulated efficiently in a quantum circuit is straightforward.
 We thus focus on the other direction.
-Suppose a certain quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step.
-Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$.
+Suppose a quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step.
+Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state right after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$.
 The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition.
@@
 As the transition rate $\lambda$ is restricted by the energy gap $\Delta$, the computation time is $\mathcal{O}(T/\Delta)$.
-IV Combinatorial optimization by Markovian quantum computation
+IV Combinatorial optimization by cooling-based quantum computation
 IV.1 Notation
-Consider combinatorial optimization problems that aim to find out the optimal solutions from a set of $2^{N}$ possible configurations represented by $N$-bit string $z\equiv z_{1}z_{2}\cdots z_{N}$.
+Consider combinatorial optimization problems that aim to find the optimal solutions from a set of $2^{N}$ possible configurations represented by $N$-bit string $z\equiv z_{1}z_{2}\cdots z_{N}$.
 The optimality of each configuration is evaluated by a non-negative cost function $E(z)$.
@@
 This problem is straightforwardly translated into the ground-state finding problem, where the cost function $E(z)$ is identified with the energy eigenvalue in Hamiltonian (1).
-Here, we set $E(z)$ to be integer multiples of a parameter $\Delta>0$.
+Here, we set $E(z)$ as integer multiples of a parameter $\Delta>0$.
+For example, for boolean satisfiability (SAT) problems, $E(z)$ is the number of unsatisfied clauses by the configuration $z$ [12].
 The cavity frequencies in Hamiltonian (2) can then be chosen as integer multiples of $\Delta$.
 By choosing $M$ to be the maximum of $E(z)$, all possible transitions in the problem Hamiltonian can be coupled to a cavity transition.
-This imposes a polynomial overhead to the computational resource.
 IV.2 Reproduction of Grover’s search algorithm
@@
 The quadratic speedup of Grover’s algorithm is proven to be optimal unless the structure of the function $f(z)$ is exploited somehow [35, 36, 37].
-To perform an analogous task in MQC, we adjust the Hamiltonian (2) as
+To perform an analogous task in CQC, we adjust the Hamiltonian (2) as
 $\begin{split}H&=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}\\ &\quad+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),\end{split}$ (5)
@@
 The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm.
-One advantage of MQC implementation is that the end of computation is signaled by a detection of the cavity photon.
+One advantage of CQC implementation is that the end of computation is signaled by a detection of the cavity photon.
 This is especially beneficial when the number of solutions $n_{0}$ is unknown.
@@
 To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level.
-The increase in the number of energy levels leads to the decrease in the degeneracy of each level, which diminishes the collective effect.
+The increase in the number of energy levels leads to a decrease in the degeneracy of each level, which diminishes the collective effect.
 Moreover, it results in an increased number of transitions required to reach the ground state.
@@
 In the previous example reproducing Grover’s search algorithm, the transition term generates an all-to-all interaction with the identical transition strength.
-This makes the problem Hamiltonian completely featureless except for each energy level having a different degeneracy.
+This makes the problem Hamiltonian utterly featureless except for each energy level having a different degeneracy.
 While the absence of any local energy minimum is advantageous, the transition strength decreases superpolynomially with the number of qubits.
@@
 As the distance between two configurations is at most $N$, any state $|z\rangle_{P}$ can reach the ground state in at most $N$ transitions in principle.
-However, we do not know a general rule to take the right path for every state.
-For this reason, we take the direction of lowering the energy, even though it does not coincide with the direction of approaching the ground state.
+However, we do not know a general rule for every state to take the right path.
+For this reason, we take the direction of lowering the energy, even though it does not necessarily coincide with the direction of approaching the ground state.
 This mismatch is the origin of local energy minima.
@@
 It is instructive to envisage this as an $N$-regular graph with $2^{N}$ vertices.
-Each vertex is endows with a potential determined by the problem Hamiltonian, and the transition occurs along the edges toward the direction of not increasing the potential.
+Each vertex is endowed with a potential determined by the problem Hamiltonian, and the transition occurs along the edges toward the direction of not increasing the potential.
 Once the population is trapped in a local potential minimum, a high-order transition is needed to get out of it.
@@
 However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition.
-Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$.
-Now taking the cavity modes into account, the last condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$.
+Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$ (simply put, the transition diagram has a mirror symmetry).
+This rule is easily verifiable with simple numerical simulations, which we do not present here.
+Now taking the cavity modes into account, the condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$.
 This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists.
 By incorporating multiple cavity modes, various tunnelling-out transition channels can appear.
-For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play a crucial role in overcoming local minima.
-To illustrated the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [30, 22]).
-To be specific, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem.
+For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play an important role in overcoming local minima.
+To illustrate the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [38, 30]).
+Specifically, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem.
 The problem Hamiltonian is constructed as follows.
 Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$.
-For a given value of $\bar{z}$, we turn each step of the calculation into an energy term in the problem Hamiltonian.
+For a given value of $\bar{z}$, we turn each calculation step into an energy term in the problem Hamiltonian.
 Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian.
-For example, the first energy term has value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{b}$ and $\Delta$ otherwise, and so on.
+For example, the first energy term has a value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has a value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{2}$ and $\Delta$ otherwise, and so on.
 The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.
-Figure 2: Numerical simulation of MQC for integer factoring of 35.
+Figure 2: Numerical simulation of CQC for integer factoring of 35.
 Dotted curves represent the energy with respect to the number of cooling cycles, averaged over $10^{3}$ samples, for (a) $\alpha_{0}=1$ and (b) $\alpha_{0}=0$.
@@
 2(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state.
-As a comparison, we have changed $a_{0}$ to zero in Fig.
-2(b) to test the effectiveness of the aforementioned mechanism for tunneling out from local minima.
-In the absence of such a mechanism, the system almost always evolves into local minima, which appear to be present in the first excited level.
+As a comparison, we have changed $a_{0}$ to zero for Fig.
+2(b) to test the effectiveness of the mechanism for tunneling out from local minima.
+Without such a mechanism, the system almost always evolves into local minima, which appear to be present at the first excited level.
 A remark is in order.
-Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears to be effective, as demonstrated in Fig.
+Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears effective, as demonstrated in Fig.
 2(a).
 However, it should be noted that the optimal choice of $\lambda$ depends on the characteristics and size of the problem.
-In general, smaller $\lambda$ is preferable as the collective effect enhance the transition rate.
+In general, smaller $\lambda$ is preferable as the collective effect enhances the transition rate.
 This effect is particularly important for combinatorial optimization problems, which are characterized by large degeneracies in the problem Hamiltonian.
-As the degeneracies of low energy levels are much smaller in general, the collective effect is also weaker in the low-energy part of the spectrum.
-This means that there exists a parametric regime of $\lambda$ where the heating is sufficiently avoided for the ground state, but not for higher energy states.
-On the other hand, by choosing smaller $\lambda$, leading to slower transitions, heating could be avoided throughout the entire energy spectrum.
-While the cooling works in both cases, it is unclear which one is more efficient.
+As the degeneracies are generally very large in the middle of the energy spectrum, the collective effect is also stronger in that region.
+Consequently, there exists a parametric regime of $\lambda$ where the heating is sufficiently avoided for the ground state, but not for higher energy states.
+On the other hand, by choosing smaller $\lambda$, leading to slower transitions, one could reach the regime where heating is avoided throughout the entire energy spectrum.
+While the cooling is feasible in both cases, it is not clear which case is more efficient.
 V Discussion
-In this work, we have introduced the notion of Markovian quantum computation, along with a few illustrative examples.
-The idea is quite straightforward: cooling the system with a problem Hamiltonian through energy-lowering transitions, as in cavity cooling.
-An essential part here is to choose an appropriate transition term that generates off-diagonal elements in the diagonal basis of the problem Hamiltonian.
-There exists a trade-off between the connectivity and strength of the transitions.
+In this work, we have introduced the notion of cooling-based quantum computation along with a few illustrative examples.
+An important part is to choose an appropriate transition term, which generates off-diagonal elements in the diagonal basis of the problem Hamiltonian.
+There exists a trade-off between the connectivity and strength of the transition.
 We have discussed two opposite limits when the ground state is the solution of a combinatorial optimization problem.
 In both cases, the running time generally increases superpolynomially with the system size.
-It is unclear if the performance can be substantially improved by interpolating between the two opposites.
-MQC can also be used to find the ground state of general quantum Hamiltonians.
-The most important requirement is that the ground state is not heated during the process to make it a steady state.
-This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap can be resolved by a proper choice of the transition term and the parameters.
-Consequently, the inverse of the spectral gap determines the basic time scale of the computation.
-In principle, a finite number of cavity modes can cover the transitions across the entire spectrum of the problem Hamiltonian, where the required number of cavity modes is given roughly by the ratio of the energy range of the problem Hamiltonian to the characteristic transition rate.
-This means that the number of cavity modes increases polynomially with the system size for gapped Hamiltonians.
-However, choosing an appropriate transition term and estimating the resulting cooling time depend heavily on the system and are analytically challenging.
+It is an open question whether the performance can be substantially improved by interpolating between the two opposites.
+In fact, the idea of CQC can be used to prepare the ground states of general quantum Hamiltonians, although we focused on the analyzable cases.
+The most essential requirement is that the ground state should not be heated to make it a steady state.
+This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap is resolved.
+Consequently, the inverse of the spectral gap determines the basic time scale of the cooling, which is a natural consequence of the energy-time uncertainty principle.
+For gapped Hamiltonians, as the transition strength can be determined finitely without the risk of heating the ground state, a finite number of cavity modes can cover the transitions across the entire spectrum in principle, where the number is roughly given by the ratio of the whole energy range to the characteristic transition rate.
+This means that the number of cavity modes increases only polynomially with the system size for gapped Hamiltonians.
+However, choosing an appropriate transition term is heavily dependent on the Hamiltonian, hence analytically challenging.
+An interesting open question is the precise relationship between CQC and AQC.
+In particular, the abundance and depth of local energy minima in CQC might be strongly correlated with the spectral gap during AQC [14, 32, 33].
+If not, the two methods can differ in efficiency for particular ground-state findings.
+In fact, the transition terms $H_{T}$ in Eqs.
+(5) and (8) are the two conventional initial Hamiltonians in AQC.
+Consequently, when Trotter-decomposed, the infinitesimal unitary evolutions of CQC with Eqs.
+(5) and (8) are identical to those of AQC, although their arrangements and the inclusion or exclusion of cavities make differences.
+This might be a clue to understanding one of the two more deeply from inspecting the other.
 Acknowledgements.

### v4

@@
-Quantum Computation by Cooling
+Markovian Quantum Computation
 Abstract
@@
 However, its use of an adiabatic evolution depending on the spectral gap of an intricate many-body Hamiltonian makes its analysis daunting.
-While it is plausible to directly cool the final gapped system of the adiabatic evolution instead, an analysis of such a scheme is surprisingly missing.
+While it is plausible to directly cool the final gapped system of the adiabatic evolution instead, the analysis of such a scheme on a general ground is missing.
 Here, we propose a specific Hamiltonian model for this purpose.
@@
 The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size.
-This case turns out to be equivalent to Grover’s quantum search algorithm, aside from some additional advantages that CQC brings.
+This case turns out to be equivalent to Grover’s quantum search algorithm, aside from some advantages that CQC brings.
 The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begin to pose challenges.
-In this case, high-order transitions, which are required to get out of local energy minima, become the bottleneck.
-These transitions require a time exponentially large in the order, which can increase with the system size in general.
+The local minima can be overcome by introducing high-order tunnel-out transitions.
 We perform the numerical simulation for a particular combinatorial optimization problem to support this idea.
@@
 Our aim is to introduce a parameter $0<\lambda\ll\Delta$ to treat $\lambda H_{T}$ as a perturbation to $H_{P}$.
-If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, where the influence of the degeneracy is ignored.
-In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition.
+If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle_{P}$ to $|z^{\prime}\rangle_{P}$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|{}_{P}\langle z^{\prime}|\lambda H_{T}|z\rangle_{P}|$, where the influence of the degeneracy is ignored.
+In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|{}_{P}\langle z^{\prime}|\lambda H_{T}|z\rangle_{P}|$ and couple it to the transition.
 Here, the subscript $m$ is the index for different cavities.
@@
 Introducing $M$ different cavity modes, our final Hamiltonian reads
-$\begin{split}H&=H_{P}+\sum_{m=1}^{M}\omega_{m}c_{m}^{\dagger}c_{m}\\ &\quad+\lambda H_{T}\otimes\left[a_{0}I+\sum_{m=1}^{M}(c_{m}+c_{m}^{\dagger})\right].\end{split}$ (2)
+$H=H_{P}+\sum_{m=1}^{M}\omega_{m}c_{m}^{\dagger}c_{m}+\lambda H_{T}\otimes\left[a_{0}I+\sum_{m=1}^{M}(c_{m}+c_{m}^{\dagger})\right].$ (2)
 Here, $a_{0}\in\{0,1\}$ determines whether the transitions preserving the cavity photon numbers are allowed.
@@
 In this case, $M$ cavity modes can be simulated with $M$ qubits.
-Initially, the system is prepared in state $2^{-N/2}\sum_{z=0}^{2^{N}-1}|z\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$.
-A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state.
-Here, the measurement of the cavity photon is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
-The computation is done by repeating this cycle.
-The required number of cycles depends on the problem and strategy.
-One possible strategy is to repeat until the cavity photon is not detected for an enough number of cycles.
-This signals with high probability that the system has reached either the ground state or local minima.
-For NP problems that are guaranteed to have at least one solution, deciding which is the case is straightforward.
-Otherwise, one may repeat the same computation to increase the confidence.
-Given that the unknown ground state is not verifiable in general, the decision is probabilistic after all.
+Our cooling protocol is a stochastic process performed by repeating the following cycle.
+To simplify the description, suppose $a_{0}=0$ and the system is initially in state $|z_{0}\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$ with arbitrary $z_{0}$.
+The energy of this state is $E(z_{0})$.
+First, we let the system evolve unitarily under Hamiltonian (2).
+This causes the system to undergo coherent oscillations between the initial state and a superposition of all states $|z_{m}\rangle_{P}|1\rangle_{m}$ satisfying $|E(z_{0})-\{E(z_{m})+\omega_{m}\}|\lesssim|{}_{P}\langle z_{m}|\lambda H_{T}|z_{0}\rangle_{P}|$, where it is understood that $|z_{m}\rangle_{P}$ may represent a superposition of multiple energy levels of $H_{P}$ in cases of degeneracy.
+After a certain period (explained below), each cavity state is measured.
+If a cavity photon, say $|1\rangle_{m}$, is detected, the system collapses into state $|z_{m}\rangle_{P}|1\rangle_{m}$, where $|z_{m}\rangle_{P}$ has a lower energy than $|z_{0}\rangle_{P}$ by $\omega_{m}$.
+In this case, we empty the cavity by resetting the state to $|0\rangle_{m}$ so that the next cooling cycle starts from the lower-energy state $|z_{m}\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$.
+If no cavity photon is detected, the system simply collapses back into the initial state of the current cycle.
+Note that in any case, the cycle ends by emptying all cavities.
+This information-discarding process simulates a zero-temperature reservoir.
+During the cycle, heating is avoided as long as $\omega_{m}\gg|{}_{P}\langle z_{m}|\lambda H_{T}|z_{0}\rangle_{P}|$ for all $m$.
+Precise control of the operation time is not essential as the cooling cycle is repeated without the risk of heating; given that the probability of detecting a cavity photon in a single cycle is finite (depending on the duration), the expected number of cycles to lower the energy is also finite.
+Having said that, the optimal duration of the unitary evolution is given by the inverse of the oscillation frequency multiplied by $\pi/2$.
+However, as this optimal time varies for each transition path and multiple oscillations take place simultaneously, finding a single optimal duration is generally challenging.
+The case for $a_{0}=1$ is elaborated in Sec.
+IV.
+The measurement of the cavity state allows us to monitor the progress of cooling.
+It is a reasonable strategy to repeat the cooling cycle until cavity photons are not detected for a sufficient number of cycles.
+This signals with high probability that the system has reached either the ground state or a local minimum.
+For NP problems that are guaranteed to have at least one solution, distinguishing between the two cases is straightforward.
+Otherwise, to increase confidence, one may repeat the computation, possibly varying the transition term $H_{T}$.
+Given that an unknown ground state is not verifiable in general, the decision is probabilistic after all.
 We will implicitly assume that the evolution by the Hamiltonian (2) is simulated on a universal quantum computer, which is reasonable given the complexity of the Hamiltonian.
 In our analysis below, the polynomial overhead associated with this implementation is not taken into account.
-III Equivalence to circuit models
+III Equivalence to the quantum circuit model
 Figure 1: The transition diagram into which a quantum circuit is converted for CQC.
@@
 Again, showing that CQC can be simulated efficiently in a quantum circuit is straightforward.
-We thus focus on the other direction.
+We thus focus on the other direction: a quantum-circuit computation can be efficiently simulated in CQC.
 Suppose a quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step.
@@
 $H_{P}$ $=$ $-\sum_{t=0}^{T}t\Delta|t\rangle_{C}\langle t|,$ (3)
-[…]
+$H_{T}$ $=$ $\sum_{t=0}^{T-1}(U_{t+1}^{P}\otimes|t+1\rangle_{C}\langle t|+\mathrm{H.c.}),$ (4)
 where the parameter $\Delta>0$ is an arbitrary energy scale, and the other parmeters are chosen as $M=1$, $\omega_{1}=\Delta$, and $\alpha_{0}=0$.
@@
 1.
-It can be seen that the average time needed to reach the ground state, which is the final state of the quantum circuit, is polynomial in $T$.
-As the transition rate $\lambda$ is restricted by the energy gap $\Delta$, the computation time is $\mathcal{O}(T/\Delta)$.
+Note that the off-diagonal element ${}_{P,C}\langle\phi_{t+1},t+1|\lambda H_{T}|\phi_{t},t\rangle_{P,C}=\lambda$ is constant for every $t$.
+The cooling occurs in the same manner as the sideband cooling [34].
+Initially, the system undergoes the Rabi oscillation between $|\phi_{0}\rangle_{P}|0\rangle_{C}|0\rangle_{1}$ and $|\phi_{1}\rangle_{P}|1\rangle_{C}|1\rangle_{1}$ with Rabi frequency $\lambda$.
+After a certain duration, the cavity state is measured.
+In case of measuring $|1\rangle_{1}$, we reset the cavity state to $|0\rangle_{1}$.
+The subsequent evolution then becomes a Rabi oscillation between $|\phi_{1}\rangle_{P}|1\rangle_{C}|0\rangle_{1}$ and $|\phi_{2}\rangle_{P}|2\rangle_{C}|1\rangle_{1}$.
+If $|0\rangle_{1}$ is measured, the same Rabi oscillation simply restarts.
+Repeating this cycle thus makes the state cascade down the energy levels.
+The optimal duration of the Rabi oscillation is $\pi/2\lambda$, for which the cavity photon is detected with unit probability.
+However, precise timing is not essential, as mentioned in the previous section.
+If the cavity photon is detected $n$ times, the current state of the system is $|\phi_{n}\rangle_{P}$.
+Consequently, the average time required to reach the final state is polynomial in $T$.
+As the Rabi frequency of the oscillation $\lambda\ll\Delta$ is constant at every cycle, the computation time is $\mathcal{O}(T/\Delta)$.
+The above Rabi oscillation relies on the conventional rotating wave approximation, which neglects terms that do not preserve the energy of the unperturbed Hamiltonian $H_{P}$.
+The omitted term in this approximation induces heating albeit with a small probability.
+However, this heating is essentially harmless because it does not take the state out of the subspace spanned by $\{|\phi_{t}\rangle_{P}|t\rangle_{C}\}$.
+Even if the heating occurs (with a small probability), the mere effect is to regress the cooling by a single step, causing the transition $|\phi_{t}\rangle_{P}|t\rangle_{C}\rightarrow|\phi_{t-1}\rangle_{P}|t-1\rangle_{C}$.
+Consequently, the overall computation time of $\mathcal{O}(T/\Delta)$ remains rigorously correct.
+As mentioned in the introduction, the equivalence of CQC to the quantum circuit model should be interpreted carefully.
+What we have shown above is that computations in one model can be converted to those in the other model up to polynomial overhead, indicating that they are computationally equivalent.
+However, it is important to note that the two models are motivated differently.
+CQC is basically a scheme to prepare a many-body ground state.
+While CQC is advantageous in cases where finding a ground state naturally accomplishes a computation, such as in combinatorial optimization, or finding a ground state itself is the objective, there is no inherent reason to prefer CQC over the quantum circuit model for ordinary quantum algorithms.
+The same argument applies to AQC as it shared the same motivation with CQC [19].
 IV Combinatorial optimization by cooling-based quantum computation
@@
 To perform an analogous task in CQC, we adjust the Hamiltonian (2) as
-$\begin{split}H&=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}\\ &\quad+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),\end{split}$ (5)
+$H=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),$ (5)
 where $I_{i}$ and $X_{i}$ denote the identity and Pauli $X$ operators acting on qubit $i$, respectively.
@@
 $|\phi_{1}\rangle_{P}=\frac{1}{\sqrt{2^{N}-n_{0}}}\sum_{f(z)\not=0}|z\rangle_{P}.$ (7)
-In this basis, the off-diagonal element becomes ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$.
-The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm.
-One advantage of CQC implementation is that the end of computation is signaled by a detection of the cavity photon.
+The cooling process is thus very simple.
+The unitary evolution of the cooling cycle is a Rabi oscillation between $|\phi_{0}\rangle_{P}|0\rangle_{1}$ and $|\phi_{1}\rangle_{P}|1\rangle_{1}$ occurring with the Rabi frequency ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$.
+By repeating the cooling cycle until the cavity photon is detected, one can obtain the solution state $|\phi_{0}\rangle_{P}$.
+As the optimal duration of the Rabi oscillation is proportional to the inverse of the Rabi frequency, the computation time is given by $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm.
+In comparison to Grover’s algorithm, one advantage of the CQC implementation is that the end of computation is signaled by the detection of the cavity photon.
 This is especially beneficial when the number of solutions $n_{0}$ is unknown.
-In this example, the problem Hamiltonian has only two energy levels corresponding to the solution states and the rest.
+In this example, the problem Hamiltonian has only two energy levels corresponding to the solution state and the rest.
 A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels.
@@
 Once the population is trapped in a local potential minimum, a high-order transition is needed to get out of it.
-To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\rightarrow|z_{1}\rangle_{P}\rightarrow\cdots|z_{n}\rangle_{P}$.
+To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\leftrightarrow|z_{1}\rangle_{P}\leftrightarrow\cdots\leftrightarrow|z_{n}\rangle_{P}$.
 For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved.
 For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$.
-The perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$.
-However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition.
-Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$ (simply put, the transition diagram has a mirror symmetry).
-This rule is easily verifiable with simple numerical simulations, which we do not present here.
-Now taking the cavity modes into account, the condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$.
+In this case, the $n$-th order perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$.
+However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ differ, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition.
+Consequently, the most prominent transitions result from cases where the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$ (see Fig.
+2 for typical cases).
+Figure 2: Demonstration of representative high-order transitions for $\Delta=10\lambda$, where $\Delta$ and $\lambda$ denote the energy gap and the Rabi frequency, respectively.
+The plots on the right show the population in $|z_{n}\rangle_{P}$ as a function of the rescaled time, starting from the initial state $|z_{0}\rangle_{P}$.
+The results for (a) $n=2,3,4,5,6$ and (b) $n=2,4,6,8,10$ are overlaid in the plots.
+The indistinguishability of the individual curves confirms the reliability of the perturbation treatment.
+Among various high-order transition channels, we demonstrate two representative cases in Fig. 2. The Hamiltonian employed for the figure is
+$H=\sum_{j=0}^{n}E(z_{j})|z_{j}\rangle_{P}\langle z_{j}|+\lambda\sum_{j=0}^{n-1}\left(|z_{j}\rangle_{P}\langle z_{j+1}|+|z_{j+1}\rangle_{P}\langle z_{j}|\right).$ (9)
+In Fig.
+2(a), we have chosen $E(z_{0})=E(z_{n})=0$, $E(z_{1})=E(z_{2})=\cdots=E(z_{n-1})=\Delta$, and $\lambda=\Delta/10$.
+In this case, the $n$-th-order perturbation theory states that the oscillation between $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ occurs with frequency $\Omega_{n}\simeq\Delta(\frac{\lambda}{\Delta})^{n}$.
+The plot on the right in Fig.
+2(a) shows the population in $|z_{n}\rangle_{P}$ with respect to the rescaled time (in unit of $\pi/\Omega_{n}$) when the initial state is $|z_{0}\rangle_{P}$.
+The plot includes the curves for $n\in\{2,3,4,5,6\}$ and they are almost indistinguishable, confirming the reliability of the perturbation treatment.
+In Fig.
+2(b), we have chosen $E(z_{j})=j\Delta$ for $j\leq n/2$ and $E(z_{j})=(n-j)\Delta$ for $j\geq n/2$.
+In this case, the oscillation frequency is given by $\Omega_{n}\simeq\Delta(\frac{\lambda}{\Delta}\frac{\lambda}{2\Delta}\cdots\frac{\lambda}{(n/2-1)\Delta})^{2}\frac{\lambda}{(n/2)\Delta}$ from the perturbation theory.
+The plot on the right shows the population in $|z_{n}\rangle_{P}$ for the initial state $|z_{0}\rangle_{P}$.
+The curves for $n\in\{2,4,6,8,10\}$ are plotted with respect to the rescaled time and they are again almost indistinguishable.
+Now, let us take the cavity modes into account.
+Note that the time evolution remains identical when any intermediate transition $|z_{i}\rangle_{P}\leftrightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\leftrightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$.
 This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists.
-By incorporating multiple cavity modes, various tunnelling-out transition channels can appear.
-For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play an important role in overcoming local minima.
-To illustrate the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [38, 30]).
+By incorporating multiple cavity modes, various tunnel-out transition channels can appear.
+For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$, equivalent to the case in Fig.
+2(a), are expected to play an important role in overcoming local minima.
+To illustrate this approach to the cooling, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [39, 30]).
 Specifically, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem.
@@
 Fig.
-2(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state.
+3(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state.
 As a comparison, we have changed $a_{0}$ to zero for Fig.
-2(b) to test the effectiveness of the mechanism for tunneling out from local minima.
+3(b) to test the effectiveness of the mechanism for tunneling out from local minima.
 Without such a mechanism, the system almost always evolves into local minima, which appear to be present at the first excited level.
@@
 Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears effective, as demonstrated in Fig.
-2(a).
+3(a).
 However, it should be noted that the optimal choice of $\lambda$ depends on the characteristics and size of the problem.
@@
 It is an open question whether the performance can be substantially improved by interpolating between the two opposites.
-In fact, the idea of CQC can be used to prepare the ground states of general quantum Hamiltonians, although we focused on the analyzable cases.
-The most essential requirement is that the ground state should not be heated to make it a steady state.
-This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap is resolved.
-Consequently, the inverse of the spectral gap determines the basic time scale of the cooling, which is a natural consequence of the energy-time uncertainty principle.
-For gapped Hamiltonians, as the transition strength can be determined finitely without the risk of heating the ground state, a finite number of cavity modes can cover the transitions across the entire spectrum in principle, where the number is roughly given by the ratio of the whole energy range to the characteristic transition rate.
-This means that the number of cavity modes increases only polynomially with the system size for gapped Hamiltonians.
-However, choosing an appropriate transition term is heavily dependent on the Hamiltonian, hence analytically challenging.
+The idea of CQC could be used to prepare the ground states of general gapped quantum Hamiltonians.
+A qualitative argument is the following.
+Recall that the transition from $|z_{1}\rangle_{P}$ to $|z_{2}\rangle_{P}$ occurs when the condition $|E(z_{2})-\{E(z_{1})-\omega_{m}\}|\lesssim|{}_{P}\langle z_{2}|\lambda H_{T}|z_{1}\rangle_{P}|$ is satisfied, where a cavity mode with frequency $\omega_{m}$ is involved.
+The key point is that a single cavity mode enables transitions with the associated energy difference up to a margin determined by the spectral linewidth $|{}_{P}\langle z_{2}|\lambda H_{T}|z_{1}\rangle_{P}|$.
+To ensure that the transition can distinguish the ground state from the excited state, this linewidth should be sufficiently smaller than the spectral gap above the ground state.
+Consequently, a finite spectral gap sets the upper bound on the transition rate, hence the time scale of the cooling.
+Crucially, a finite spectral gap in turn allows to have a finite linewidth for transitions.
+The entire spectral range of the problem Hamiltonian, which is $\mathcal{O}(N)$, can then be divided into $\mathcal{O}(N/\lambda)$ sectors.
+This means that a polynomial number of cavity modes suffice to cover the entire spectrum, provided that one can design appropriate transition terms $H_{T}$ along with the cavity modes to establish transition channels to the ground state for all the energy levels.
+Of course, the practical application of this idea would heavily depend on the specific Hamiltonian and hence analytically challenging.
+The requirement of a finite spectral gap is a natural consequence of the energy-time uncertainty principle.
 An interesting open question is the precise relationship between CQC and AQC.